The area og a rectangle is 36. IF the length is 9, what is the perimeter?
The answer is 26, but how do you get the answer. Please show your work.

Answer by queenofit(120)

First you must have the formula for perimeter of a rectangle
P=2L + 2W
you have the length and the perimeter values, so plug them into the formula
26 = 2(9)+2W
26 = 18 + 2w
26-18 = 2w
8= 2w
4=w
check your work by plugging the numbers into the formula
26= 2(9) + 2(4)
26=18 + 8
26=26